OpenStudy (anonymous):

A number cube with faces labeled 1–6 is rolled 90 times. About how many times would the result be greater than 2? A. 30 B. 45 C. 60 D. 75

OpenStudy (anonymous):

P(no.>2)=4/6=2/3 \[this~ is~ probability~ for~ one` throw ,multiply~ by~ 90~ and~ find~ the~ solution.\]